# Context-Aware Code Generation Framework for Code Repositories: Local, Global, and Third-Party Library Awareness

### 18 Dec 2023 v2

@@
Abstract.
Code generation tools are essential to help developers in the software development process. Existing tools often disconnect with the working context, i.e., the code repository, causing the generated code to be not similar to human developers. In this paper, we propose a novel code generation framework, dubbed $A^{3}$-CodGen, to harness information within the code repository to generate code with fewer logical errors, code redundancy, and library-related compatibility issues. We identify three categories of representative information for the code repository: local-aware information from current code file, global-aware information from other code files, and third-party-library information. Results demonstrate that by adopting the $A^{3}$-CodGen framework, we successfully extract, fuse, and feed code repository information into the LLM, generating more accurate, efficient, and highly reusable code. The effectiveness of our framework is further underscored by generating code with a higher reuse rate, compared to human developers. This research contributes significantly to the field of code generation, providing developers with a more powerful tool to address the evolving demands in software development in practice.
Code Generation, Code Evolution, Code Repository Knowledge Mining
copyright: none
1. Introduction
@@
A significant gap still persists between the code generated by LLMs and that crafted by human developers. Typically, when adding new functionality or features to a specific module, developers already had sufficient knowledge about the current task and the entire code repository. Also, according to a previous study [11], developers often expect that code generation tools could be aware of more context/knowledge by using specific frameworks/libraries [12]. In addition, they hope the generated code maximum utilizing specific third-party libraries (e.g., invocation of an API in a library) [13, 14, 15]. Above all, in contrast to the straightforward LLM-based code generation, developers especially take into account three primary categories of information:
@@
Local Information: This category encompasses information within the current working context, including defined function signatures and variables. It also covers the relative file paths of the module within the project, such as the module’s Fully Qualified Names (FQNs), like unstructured.documents.html. local information is crucial to ensuring the generated code snippet utilizes accurate variables and operates harmoniously with the other functions in the project.
@@
Global Information: This category involves functions from other code files within the same repository, accessible via fully qualified import API names. A previous research highlighted that only around 30% of methods operate independently in open-source projects [16]. Hence, these external functions are essential for implementing new functions without duplicating existing code. For example, developers import functions “from unstructured.partition.text_type import is_bulleted_text” from other modules, instead of writing a function to judge “bulleted text” from scratch.
@@
Third-Party-Library Information: This category includes third-party libraries that are already pre-installed in the code repository’s environment. These libraries can be accessed directly through import statements, enhancing the functionality without additional dependencies.
Unfortunately, none of existing code generation methods take above three information into consideration simultaneously. As the most common code generation tool, ChatGPT [8] showcases a commendable level of coding ability [17], but it solely depends on the users’ input. GitHub Copilot [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository. When these methods disregard local and global module information, they risk omitting crucial logic during code generation [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as “ModuleNotFoundError: No module named xxx.
@@
Our experiments are driven by two primary objectives: 1) To investigate the effectiveness of three distinct types of code repository information in assisting the model with code generation for specific code repositories. 2) To assess whether our framework can efficiently extract and utilize these three types of information. We crawl and filter 29 representative Python code repositories from PyPI and randomly sample 383 representative functions with comments to create our benchmark dataset. Our experimental results confirm the efficacy of incorporating three types of code repository-awareness knowledge in aiding the foundational model to reuse code repository knowledge. The $A^{3}$-CodeGen framework excels in recognizing and utilizing these three types of code repository knowledge, leading to exceptional performance metrics in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding code reuse capabilities, our framework is capable of generating concise code that seamlessly integrates with the code repository’s working environment. Furthermore, we identify optimal configurations for local and global information, enabling the model to effectively leverage essential knowledge while avoiding the potential pitfalls of information overload.
@@
  • We constructed a new benchmark dataset, namely RepoEval, to reflect the effectiveness on code-repository-aware code generation. Results demonstrate the validity of the three types of code repository information and show that our framework can effectively mine and utilize this knowledge.
2. Motivation
Figure 1. Knowledge-Starved Code Generation vs. Code Repository Aware Code Generation.
In this section, we discuss the motivation of this study.
2.1. Limitation Of Existing Code Generation Approaches
Benefit by the rapid development of Transformer structure and pre-trained LLMs in the early 2020s, code-LLMs also thrived significantly. Encoder-decoder Transformer code models were dominant, such as the milestone work CodeT5 [19]. Also, other code models like CodeBERT [20] adapt the original BERT architecture to code generation tasks, seamlessly bridging natural language understanding with code comprehension. Building on this, GraphCodeBERT [21] extend its capability by incorporating the graph structure of code to enhance performance. Whilst, more recently, the decoder-only Transformer models have become the mainstream of generative models. Codex [22] and CodeGen [23] are representatives, and many others, like [24, 25, 26, 27, 28], also belong to this category.
Nowadays, to Software Engineering developers, the LLM-empowered code generation has become an irreplaceable assistant tool in their development cycle [29, 30]. Although the code-LLM has been developed for years, those code generation tools still play a supporting role in the actual SE development. Essentially, there are significant differences between the code generated by LLMs and that crafted by human developers. Specifically, we conclude this phenomenon into the following three aspects:
@@ 2.1. Limitation Of Existing Code Generation Approaches
    Discussion 1: Providing longer prompts with richer content seems a plausible solution. With the growth of PLM size, the maximum input length increases from 512 tokens for the early models [31] to 32,000 tokens for the extended GPT-4 models [32]. However, it is still impossible to feed the entire code repository and all other relevant information into the LLMs. Furthermore, studies have demonstrated that irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performance [33]. Therefore, to maintain the practicability and effectiveness simultaneously for code generation tasks, external context information should be extracted and selectively fed into LLMs.
  • Observation 2: Existing methods focus on generating independent and often small-scale code units [34, 16]. According to a previous study [34], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (Section 7)), newly generated code still remains questionable.
@@
    Discussion 3: Similarly, instead of generating complex code snippets from scratch, importing functions that are already implemented can mitigate the shortcomings of advanced-level programming challenges. For example, Python developers are likely to import functions from networkx [^1] https://networkx.org/documentation/stable/index.html package to build a graph, not building graphs by the dictionary-like data structure.
2.2. Code Generation in Code Repository
During software development, programmers typically work within a specific code environment, which is essentially a designated code repository. This code repository encompasses all the project’s modules, functions already present in these modules, and third-party libraries. It serves as the foundational framework upon which developers build and expand their software applications. In this context, developers (especially experienced ones) must navigate and understand this code repository environment thoroughly to ensure seamless integration of newly added code with the existing one. When it comes to LLM-based code generation techniques, solely relying on inputting developer requirements to these models fails to fully utilize the potential of LLMs in interacting with existing code repository environments. Specifically, those techniques are blind to the repository’s unique naming conventions and programming styles when the completion target is relevant to cross-file context [36, 37, 38]. Therefore, instead of using a single code file, well-designed programs emphasize separating complicated functionality into independent modules, each implementing only one aspect of the desired functionality [39, 40].
@@ 2.2. Code Generation in Code Repository
Furthermore, since the model is unaware of the existence of the ”check text for bullet characters” function in the global module, it defines a new function is_bullet_text (highlighted in the blue box) in the generated code. A large part of this code duplicates the function is_bullet_text in the global module unstructured/partition/text_type.py, resulting in code redundancy. It also brings additional concerns about the correctness of the newly generated function is_bullet_text.
Additionally, since the model is unaware of which third-party libraries are available, it arbitrarily imports and uses the bs4 (BeautifulSoup4222https://pypi.org/project/beautifulsoup4/) library. This can result in the generated code that refers to a library that is not available, leading to the bug “ModuleNotFoundError: No module named ***”.
These knowledge gaps highlight the limitations of existing code generation approaches, emphasizing the need for more context-aware methods that can effectively interact with the code repository environment. In summary, a lack of knowledge about local and global modules can result in models that generate code that lacks functional constraints or contains redundancies; and a lack of knowledge about third-party libraries in the code repository can lead to models using arbitrary third-party libraries, which can cause problems with the generated code not working. As a result, generating code in specific modules of the code repository is more than just meeting functional requirements. To generate code similar to human developers, the approach must fully understand and effectively utilize this diverse contextual knowledge to ensure that newly introduced code is seamlessly coordinated with the existing code repository.
2.3. Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
@@
Figure 2. $A^{3}$-CodGen Framework.
3. $A^{3}$odGen Framework
In this section, we specifically elaborate on our novel $A^{3}$-CodGen Framework to harness the full potential of code repository knowledge for assisting LLMs in code generation.
@@ 3. $A^{3}$odGen Framework
As depicted in Fig. 2, the $A^{3}$-CodGen Framework comprises three essential phases. First, we obtain the function base and third-party library base in the code repository knowledge construction phase (Section 3.1), and they provide sufficient knowledge to support subsequent knowledge retrieval. Then, we retrieve the code repository knowledge from other code files, the third-party library repository, and the current file. After that, we organize them into 3 types of code repository aware knowledge (Section 3.2). Finally, in the knowledge utilization phase (Section 3.3), the three code repository-aware knowledge are fused into a prompt, which serves as an input to the foundation LLM, to ultimately generate codes similar to human developers with three essential awareness.
3.1. Repository Knowledge Base Construction
@@
Code repositories contain a wealth of knowledge to exploit, but this knowledge usually exists in an unstructured form and cannot be directly utilized effectively, scattering in the code repository. Therefore, we need to perform systematic knowledge extraction to transform and compress this knowledge into a structured form in order to build a knowledge base to support subsequent knowledge utilization, such as knowledge retrieval. Thus, we perform code repository information extraction (Section 3.1.1) and third-party library extraction (Section 3.1.2) to obtain the function base and third-party-library base, respectively.
Figure 3. A Piece of Code in the Module & Function Summarizer Unit
3.1.1. Code Repository Information Extraction
@@
Given a code repository, we can effortlessly obtain all source code files in the repository’s root folder. Then, for each source code file, we can extract all functions included. For each function, we perform both direct code information collection and semantic code knowledge extraction. In total, by snowballing this process, all essential repository information is extracted.
@@
Basic Code Knowledge Extraction. Abstract Syntax Tree (AST) is a tree representation of the source code, and studies have demonstrated its usefulness on various SE tasks [41]. In this study, we employ the Python library ast [^3] https://docs.python.org/3/library/ast.html to analyze the AST for each source code file in the code repository. From the AST, we extract six essential types of basic information: 1) Comment. For example, in Fig 3, the comment for the function element (circled in the yellow box) is Get all elements from pages in sequential order. 2) Signature. This knowledge illustrates the parameters and returns of a function. For example, in Fig 3, the signature of the function element (circled in the yellow box) is def elements(self) -¿ List[Element]. 3) Function Source code. 4) Class associated with the function. For example, the function element in Fig. 3 corresponds to the class Document. 5) Is_Empty tag. True means empty function, False means empty label. For example, function element has is_Empty value False. 6) The file path of the function. This reflects the relative location of the current function in the entire code repository. For example, the file path of the function element is unstructured/documents/base.py.
@@
Semantic Code Knowledge Extraction. AST is criticized for its prone to the long-term dependency problem. Also, simply employing AST potential weakens the capability to capture real and complex semantics [41]. Thus, we further maneuver Semantic Code Knowledge Extraction to enhance the basic code knowledge with additional semantic information. During this process, we extract three types of semantic code knowledge, as follows:
1) Fully Qualified Name (FQN) of Functions. FQN refers to the complete name of a function, typically composed of the module path, class name (if the function belongs to a class), and the function name. For example, in Fig. 3, the FQN of the function element (highlighted within the yellow box) is unstructured.documents.base.Document.elements. FQN is often used to represent structural information about a function, indicating which module it originates from in the code repository and, if applicable, which class within the module. Developers can import global functions from the global module using their FQNs.
2) Function Summary. Not every function is elaborated by comments. For instance, in Fig. 3 (a), the function __init__ (encircled in a red box) lacks a comment. Even though the comments may exist, their quality varies distinctly. Some do not clearly state the basic functionality and usage of the function; whilst other comments may cover too many non-essential details such as notifications, technical details, or other secondary information unrelated to the overall functionality. Over-length textual descriptions may distract people from understanding the actual usage and operational behavior of the function. Therefore, we generate the function summary to precisely describe the purpose and functionality. Much previous work has shown that LLMs are capable of in-context learning [42, 43, 44], just input a task description and some examples can make LLMs behave well. Thus we use GPT-3.5-Turbo as the foundation model to implement the function summarization task.
@@
Function Summarizer Unit. To generate natural language semantic information for a function, we design a prompt as the core of the Function Summarizer unit. This prompt creates summaries for each function. As depicted in Fig. 3, our Function Summarizer prompt consists of a task description (highlighted in purple), five demonstration examples, and the input template. The input to this unit is the function’s code (highlighted in green), and the output is the model-generated summarization for that function (highlighted in red). For instance, when provided with the code of the __init__ function within the red box in Fig. 3, the resulting summary is Initialize the object with two optional attributes: _pages and _elements.
@@
3) Vector Representations. In the previous two steps, we underlined explictly select information. In order to facilitate subsequent retrieval of global functions, it is crucial to possess an effective code representation that accurately captures the semantics of functions. Therefore, we employ a dual-pronged approach, conducting separate embeddings for both the source code and the function summaries. This results in the creation of code vector and summary vector, respectively.
3.1.2. Third-Party Library Extraction

### 03 Mar 2024 v3

@@
Abstract
Code generation tools are essential to help developers in the software development process. Existing tools often disconnect with the working context, i.e., the code repository, causing the generated code to be not similar to human developers. In this paper, we propose a novel code generation framework, dubbed $A^{3}$-CodGen, to harness information within the code repository to generate code with fewer potential logical errors, code redundancy, and library-induced compatibility issues. We identify three categories of representative information for the code repository: local-aware information from current code file, global-aware information from other code files, and third-party-library information. Results demonstrate that by adopting the $A^{3}$-CodGen framework, we successfully extract, fuse, and feed code repository information into the LLM, generating more accurate, efficient, and highly reusable code. The effectiveness of our framework is further underscored by generating code with a higher reuse rate, compared to human developers. This research contributes significantly to the field of code generation, providing developers with a more powerful tool to address the evolving demands in software development in practice.
Index Terms:
Code Reuse, Code Generation, Code Repository Knowledge Mining, Prompt Engineering
I Introduction
@@
A significant gap still persists between the code generated by LLMs and that crafted by human developers. When adding new functionality to a specific module, developers already had sufficient knowledge about the current task and the entire code repository. Also, according to a previous study [11], developers often expect that code generation tools could be aware of more context/knowledge by using specific frameworks/libraries [12]. In addition, they hope the generated code maximum utilizes specific third-party libraries (e.g., invocation of an API in a library) [13, 14, 15]. Above all, in contrast to straightforward LLM-based code generation, developers especially consider three primary categories of information:
@@
Local Information: This category encompasses information within the current working context, including defined function signatures and variables. It also covers the relative file paths of the module within the project, such as the module’s Fully Qualified Names (FQNs), like unstructured.documents.html. Local information is crucial to ensuring the generated code utilizes accurate variables and operates harmoniously with the other functions in the project.
@@
Global Information: This category involves functions from other code files within the same repository, accessible via fully qualified import API names. A previous research highlighted that only around 30% of methods operate independently in open-source projects [16]. Hence, these external functions are essential for implementing new functions without duplicating code. For example, developers import functions “from unstructured.partition.text_type import is_bulleted_text” from other modules, instead of writing a function to judge “bulleted text” from scratch.
@@
Third-Party-Library Information: This category includes third-party libraries pre-installed in the code repository’s environment. These libraries can be accessed directly through import statements, enhancing functionality without extra dependencies.
Unfortunately, none of existing code generation methods take above three information into consideration simultaneously. As the most common code generation tool, ChatGPT [8] demonstrates commendable coding ability [17] but relies solely on user input. GitHub Copilot [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository [^1] https://githubnext.com/projects/copilot-view/. When these methods disregard local and global module information, they risk omitting crucial logic during code generation [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as “ModuleNotFoundError: No module named xxx.
@@
Our experiments are driven by two primary objectives: 1) To investigate the effectiveness of three distinct types of code repository information in assisting the model with code generation for specific code repositories. 2) To assess whether our framework can efficiently extract and utilize these three types of information. We crawl and filter 29 representative Python code repositories from PyPI, assembling 13,784 functions. To ensure a diverse benchmark dataset, we conduct a sampling process, selecting representative functions from each repository to total 383 functions. We subsequently label these benchmarks and the 3,830 model outputs generated under different configurations to evaluate the performance of $A^{3}$-CodGen. Our experimental results confirm the efficacy of incorporating three types of code repository-awareness knowledge in aiding the foundational model to reuse code repository knowledge. The $A^{3}$-CodeGen framework excels in recognizing and utilizing these three types of code repository knowledge, leading to exceptional performance metrics in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding code reuse capabilities, our framework is capable of generating concise code that seamlessly integrates with the code repository’s working environment. Furthermore, we identify optimal configurations for local and global information, enabling the model to effectively leverage essential knowledge while avoiding the potential pitfalls of information overload.
@@
  • We construct a new benchmark dataset, namely RepoEval, to reflect the effectiveness on code-repository-aware code generation. Results demonstrate the validity of the three types of code repository information and show that our framework can effectively mine and utilize this knowledge.
  • We open-source our replication package [19], including the dataset, source code, and experimental results, for follow-up research.
II Motivation
Figure 1: Knowledge-Starved Code Generation vs. Code Repository Aware Code Generation.
II-A Limitation of Existing Code Generation Approaches
Benefiting from the rapid development of Transformer structures and pre-trained LLMs in the early 2020s, code-LLMs experienced significant growth. Transformer code models, exemplified by milestone works such as CodeT5 [20], CodeBERT [21], and GraphCodeBERT [22], dominated the scene. More recently, decoder-only Transformer models have become the mainstream in generative models. Represented by Codex [23], CodeGen [24], and others like [25, 26, 27, 28, 29].
Nowadays, to Software Engineering developers, LLM-empowered code generation has become an irreplaceable assistant tool in their development cycle [30, 31]. Although the code-LLM has been developed for years, those code generation tools still play a supporting role in the actual SE development. Essentially, there are significant differences between the code generated by LLMs and that crafted by human developers. Specifically, we conclude the phenomenon in the following threefold:
@@ II-A Limitation of Existing Code Generation Approaches
    Discussion 1: Providing longer prompts with richer content seems a plausible solution. With the growth of LLM size, the maximum input length increases from 512 tokens for the early models [32] to 32,000 tokens for the extended GPT-4 models [33]. However, it is still impossible to feed the entire code repository and all other relevant information into the LLMs. Furthermore, studies have demonstrated that irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performance [34]. Therefore, to maintain the practicability and effectiveness simultaneously for code generation tasks, external context information should be extracted and selectively fed into LLMs.
  • Observation 2: Existing methods focus on generating independent and often small-scale code units [35, 16]. According to a previous study [35], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (Section VII)), newly generated code still remains questionable.
@@
    Discussion 3: Similarly, instead of generating complex code snippets from scratch, importing functions that are already implemented can mitigate the shortcomings of advanced-level programming challenges. For example, Python developers are likely to import functions from networkx package to build a graph, not building graphs by the dictionary-like data structure.
II-B Code Generation in Code Repository
During software development, programmers typically work within a specific code environment, essentially a designated code repository. This repository encompasses all project modules, functions in these modules, and third-party libraries, serving as the foundational framework for building and expanding software applications. In this context, developers, especially experienced ones, must thoroughly navigate and understand this code repository environment to ensure seamless integration of newly added code with the existing one. When it comes to LLM-based code generation techniques, solely relying on inputting developer requirements to these models fails to fully utilize the potential of LLMs in interacting with existing code repository environments. Specifically, these techniques are blind to the repository’s unique naming conventions and programming styles, particularly when the completion target is relevant to cross-file context [37, 38, 39]. Instead of using a single code file, well-designed programs emphasize separating complicated functionality into independent modules, each implementing only one aspect of the desired functionality [40, 41].
@@ II-B Code Generation in Code Repository
Furthermore, since the model is unaware of the existence of the “check text for bullet characters” function in the global module, it defines a new function is_bullet_text (highlighted in blue) in the generated code. A large part of this code duplicates the function is_bullet_text in the global module unstructured/partition/text_type.py, resulting in code redundancy. It also brings additional concerns about the correctness of the newly generated function.
Additionally, since the model is unaware of which third-party libraries are available, it arbitrarily imports and uses the bs4 (BeautifulSoup4) library. This can result in the generated code that refers to a library that is not available, leading to the bug “ModuleNotFoundError: No module named ***”.
These knowledge gaps highlight the limitations of existing code generation approaches, highlighting the need for more context-aware methods that can effectively interact with the code repository environment. In summary, a lack of knowledge about local and global modules can result in models generating code lacking functional constraints or containing redundancies. A lack of knowledge about third-party libraries in the code repository can lead to models using arbitrary third-party libraries, causing problems with the generated code not working. As a result, generating code in specific modules of the code repository is more than just meeting functional requirements. To generate code similar to human developers, the approach must fully understand and effectively utilize this diverse contextual knowledge to ensure that newly introduced code seamlessly coordinates with the existing code repository.
II-C Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
@@
Figure 2: $A^{3}$-CodGen Framework.
III $A^{3}$odGen Framework
@@
In this section, we elaborate on our novel $A^{3}$-CodGen Framework to harness the full potential of code repository knowledge for assisting LLMs in code generation. As depicted in Fig. 2, the $A^{3}$-CodGen Framework comprises three essential phases. First, we obtain the function base and third-party library base (Section III-A), providing sufficient knowledge to support subsequent retrieval. Then, we retrieve the code repository knowledge from other code files, the third-party library repository, and the current file. After that, we organize them into 3 types of code repository aware knowledge (Section III-B). Finally, the three code repository-aware knowledge are fused into a prompt (Section III-C), serving as input to the foundation LLM, to ultimately generate codes with three essential awareness.
III-A Repository Knowledge Base Construction
@@
Code repositories contain a wealth of knowledge to exploit, but this knowledge usually exists in an unstructured form and cannot be directly utilized effectively, scattering in the code repository. Therefore, we need to perform systematic knowledge extraction to transform and compress this knowledge into a structured form in order to build a knowledge base to support subsequent knowledge utilization, such as knowledge retrieval. Thus, we perform code repository information extraction (Section III-A1) and third-party library extraction (Section III-A2) to obtain the function base and third-party-library base, respectively.
III-A1 Code Repository Information Extraction
@@
For a code repository, we can effortlessly obtain all code files in the repository’s root folder. Then, for each file, we extract all included functions, gathering both basic code information and semantic code knowledge. This process snowballs to extract all essential repository information.
@@
Basic Code Knowledge Extraction. Abstract Syntax Tree (AST) is a tree representation of the source code, and studies have demonstrated its usefulness on various SE tasks [42]. In this study, we employ the Python library ast to analyze the AST for each source code file in the code repository. From the AST, we extract six essential types of basic information: 1) Comment. 2) Signature. This knowledge illustrates the parameters and returns of a function. For example, the signature of the function _construct_text is def _construct_text(tag_elem, include_tail_text)-$>$str. 3) Function Source code. 4) Class associated with the function (if a function is a member of a class). 5) Is_Empty tag. True means empty function, False means empty label. For example, function _construct_text has is_Empty value False. 6) The file path of the function. This reflects the relative location of the current function in the entire code repository.
@@
Semantic Code Knowledge Extraction. AST is criticized for its prone to long-term dependency problems. Also, simply employing AST potential weakens the capability to capture real and complex semantics [42]. Thus, we extract three types of semantic code knowledge to enhance the basic code knowledge, as follows:
1) Fully Qualified Name (FQN) of Functions. FQN indicates the function’s origin in the code repository and, if applicable, the specific class within the module. For example, in the local module of the Fig. 1, the function _construct_text (highlighted in yellow) has the FQN unstructured.documents.html._construct_text.
2) Function Summary. Not every function is elaborated by comments. Even though the comments may exist, their quality varies distinctly. Therefore, we generate the function summary to precisely describe the purpose and functionality.
@@
Function Summarizer Unit. To generate natural language semantic information for a function, we design a prompt as the core of the Function Summarizer unit. This prompt creates summaries for each function. Our Function Summarizer prompt consists of a task description, five demonstration examples, and the input template. The input to this unit is the function’s code, and the output is the model-generated summarization for that function.
@@
3) Vector Representations. In the previous two steps, we underlined explictly select information. To ensure efficient retrieval of global functions later on, having an effective code representation that faithfully captures function semantics is essential. Hence, we employ a dual-pronged approach, conducting separate embeddings for both the source code and the function summaries. This results in the creation of code vector and summary vector, respectively.
III-A2 Third-Party Library Extraction